given market demand Q=100-P if the market supply function for the smaller firms is given by S=0.5P. And the cost function of the dominant firms is TC=10+40QD, where QD= output of the dominant firm. a) Find the market price and output of the dominant firm at equilibrium. b) Find the output level to be supplied by the smaller firms.


1
Expert's answer
2021-04-05T07:22:44-0400

a) The market price and output of the dominant firm at equilibrium are:

MR = TR'(Q) = 100 - 2Q,

MC = TC'(QD) = 40,

MR = MC,

100 - 2Q = 40,

Q = 30 units,

P = 100 - 30 = 70.

b) The output level to be supplied by the smaller firms is:

Q = S,

100 - P = 0.5P,

P = 66.67,

Q = 33.33 units.